OpenStudy (anonymous):

cos4L+4cos2L+3=8cos^4L

OpenStudy (***[isuru]***):

u can write \[\cos \ 2x = 2\cos^{2} x - 1\] use this and expand ur expression until u get the answer

OpenStudy (jdoe0001):

\(\bf cos^4(L)+4cos^2(L)+3=8cos^4(L)\quad ?\)

OpenStudy (***[isuru]***):

it;s \[\cos( 4L) + 4\cos(2L) = 8\cos^{4}L\]

OpenStudy (***[isuru]***):

use... \[\cos(2L) = 2\cos^{2}L -1\] and exapand \[4 \cos (2L)...\] it will be something like this \[4 [ 2\cos^{2}L -1 ]\]

OpenStudy (***[isuru]***):

then... take \[\cos (4L) \ as \ \cos(2\times2L)\] then u can write.. \[\cos(4L) = \cos(2 \times2L) =2\cos^{2}2L - 1\]

OpenStudy (***[isuru]***):

so u can write... \[\cos (4L) + 4\cos(2L) + 3\] as \[2\cos^{2}(2L) - 1 + 4[2\cos^{2}L - 1] + 3\]

OpenStudy (***[isuru]***):

now expand the \[2\cos^{2} (2L) \] using \[\cos(2L) =2\cos^{2}L - 1\] where u will get \[2[ 2\cos^{2} L - 1]^{2}\] now expand this expression.. \[2( 2\cos^{2}L - 1)^{2} = 2[4\cos^{4}L + 1- 4\cos^{2} L] = 8\cos^{4} L+ 2-8\cos^{2}L\]

OpenStudy (***[isuru]***):

now just write all the parts together... u will get ur answer.. \[8\cos^{4}L+2 -8\cos^{2}L -1 +4(2\cos^{2}L - 1) + 3\]

OpenStudy (***[isuru]***):

simplify this one( i mean remove brackets then add and subtract ...) u will get ur answer ...:) got it :) ?